OpenStudy (anonymous):

Minimize S = x + 14y with xy = 14 and both x and y > 0. Minimize S = x + 14y with xy = 14 and both x and y > 0. @Mathematics

OpenStudy (anonymous):

xy = 14 x = 14/y 14/y + 14y = S 14/y + 14y^2/y = S (14+14y^2)/y = S f'(S) = 14-14/y^2 = 0 y = -1,1 y>0 so y =1 x = 14

OpenStudy (anonymous):

You are the BEST !

OpenStudy (anonymous):

why did you raise y^2 boss?

OpenStudy (anonymous):

This derivative invloves the quotient rule

OpenStudy (anonymous):

14/y + 14y = S 14/y + 14y^2/y = S <- this step beginning to solve for f'(x) correct?

OpenStudy (anonymous):

GOT IT ! THANKS :)